OpenStudy (anonymous):

A company has determined that when x items are sold daily, the profit P is given by P(x)=-0.002x^2+4.3x-120 What is the profit if 700 items are made? How many items should be made daily in order to maximize the companys profit? I know the first one is 1,910, but how do I maximize the company's profit?

OpenStudy (anonymous):

This basically asks you to find the vertex of the equation

OpenStudy (anonymous):

find the derivative and set it to zero

OpenStudy (anonymous):

solve for x then

OpenStudy (anonymous):

What formula would I use?

OpenStudy (anonymous):

I don't know any formulas but I do know how to find it using a derivative

OpenStudy (anonymous):

Ok, any advice on how I should proceed?

OpenStudy (anonymous):

Find the derivative

OpenStudy (anonymous):

So if I get a derivative of 4.3-0.004x, what should I do then?

OpenStudy (anonymous):

set it to zero and solve for x

OpenStudy (anonymous):

4.3-0.004x=0

OpenStudy (anonymous):

4.3-0.004x=0 -0.004x=-4.3 x=-4.3/-0.004 x=1075

OpenStudy (anonymous):

1075 items

OpenStudy (anonymous):

i see, so the derivative is the key, awesome, thanks!

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=-0.002x%5E2%2B4.3x-120 if you graph it you can see the highest point right? That point is called the vertex and you can find it using the derivative

OpenStudy (anonymous):

Ah, ok I got you.